OpenStudy (anonymous):

A marble is drawn at random from the bag shown below. The bag contains 3 green, 4 yellow, 5 blue, and 8 pink marbles. What is P(not yellow)?

OpenStudy (anonymous):

total possible outcomes = 3 + 4 + 5 + 8 = 20 Not yellow means green, blue or pink balls Not yellow balls = 3 + 5 + 8 = 16 P(not yellow) = 16/20 = 4/5